### v1

@@ 6.2 Influence of boundary conditions on the long-term evolution
Thermal boundary conditions play a critical role in determining cooling timescales, as they govern energy losses through surface photon emission. Notably, significant differences arise when comparing non-magnetized envelope models (Gudmundsson et al, 1983; Potekhin et al, 1997) with magnetized ones (Potekhin et al, 2003, 2015b), for both light-element (hydrogen) and heavy-element (iron-like) compositions. Light-element envelopes, commonly used for accreting sources, produce luminosities up to an order of magnitude higher than heavy-element envelopes during the neutrino-cooling phase. Due to the efficient surface photon losses, light-element envelope models cool down faster once the photon-cooling era begins, rendering the objects hardly detectable in terms of thermal $X$-rays ($L_{X}\lesssim 10^{32}$ erg/s), much before than the heavy-element models. This effect is further amplified by strong magnetic fields, as shown by Dehman et al (2023a). They showed that the predicted thermal X-ray luminosity varies significantly based on multiple factors, particularly the envelope’s properties (see Sect. 2.2.5 for details). Their findings are summarized in Fig. 21, which compares outcomes under different assumptions about the magnetic field, composition (iron versus light elements like hydrogen, typical in accreted envelopes), or the internal current distribution (core-threading field versus crustal-confined field).
@@
Figure 21: Luminosity curves for four envelope models: non-magnetised heavy envelopes (solid, Gudmundsson et al 1983), non-magnetised light envelopes (dots, Potekhin et al 1997), magnetised light envelope (dashes, Potekhin et al 2003), and magnetised heavy envelope (dot-dashes, Potekhin et al 2015b). Left panels show models with crust-confined magnetic fields, while right panels show those with core-dominated magnetic fields, in all cases consisting of initial large-scale components only. Results are presented for two initial polar field strengths: $B=5\times 10^{14}$ G (top panels) and $B=10^{13}$ G (bottom panels). Figure courtesy of Dehman et al (2023a).
@@
Beyond initial conditions, early-time ($t\lesssim 100$ yr) field reshaping and boundary-condition effects, the key question already mentioned in the previous section is arguably the location of the electric currents sustaining the magnetic field. Crustal-confined fields have been extensively studied and many works (Pons and Geppert, 2007; Viganò et al, 2012, 2013; Gourgouliatos et al, 2013; Gourgouliatos and Cumming, 2014b, a; Viganò et al, 2021; Gourgouliatos and Pons, 2019; De Grandis et al, 2020, 2022; Dehman et al, 2023c, b) generally agree on the overall picture of the Hall-driven dynamics in these configurations.
For typical field strengths of $10^{14}$ G, and starting from a predominantly poloidal dipolar field, we observe a stage dominated by the Hall drift (readjusting from initial conditions), which creates higher-order multipoles, followed by a quasi-stationary Ohmic stage. This structure, which has been called the Hall attractor (Gourgouliatos and Cumming, 2014a; Bransgrove et al, 2018), is characterized by a nearly constant angular velocity of the ”electron” fluid ($\Omega\approx j/en_{e}r$) along each poloidal field line, and proportional to the magnetic flux. It is worth noting that Hall drift can significantly accelerate magnetic field dissipation by steadily channeling magnetic energy to smaller scales, where Ohmic dissipation is more efficient.
@@
As an example, in Fig. 22 we show three snapshots of the evolution of a simple crustal-confined axisymmetric model, initially a $\ell=1$ poloidal field with $B_{p}=10^{14}$ G (labeled as model A14 in Viganò et al 2013). Such very simple initial configuration allows one to analyse and capture several general basic features characterizing the Hall-dominated dynamics. Let us recap the most important facts:
@@
    At this stage, most of the current flows near the crust/core interface, where magnetic energy dissipation is governed by the resistivity of this region. In this particular model, a highly resistive layer in the nuclear pasta region causes rapid magnetic field decay, directly affecting the observable rotational properties of X-ray pulsars (Pons et al, 2013).
@@
Figure 23: Magnetic energy in the crust (normalized to the initial value) as a function of time, for the same model of Fig. 22. The solid lines correspond respectively to the total magnetic energy (black), the energy in the poloidal component (red), and the energy in the toroidal component (blue). The dashed line shows the evolution of the same model when the Hall term is deactivated (only Ohmic dissipation). Figure courtesy of Viganò et al (2013).
@@
In 3D, details become even more relevant. Fig. 24 highlights the role of complex, multipolar magnetic structures close to the star surface in producing anisotropies in the temperature evolution. Such anisotropies have direct implications for the observable surface emission, potentially biasing cooling-age estimates and the interpretation of $X$-ray spectra, if oversimplified. This underscores the need for careful multi-dimensional treatments of the heat diffusion equation, rather than relying on one-dimensional cooling models (De Grandis et al, 2021; Igoshev et al, 2021, 2023; Dehman et al, 2023b; Ascenzi et al, 2024).
@@
Figure 24: 3D visualizations of different magnetic field configurations and their impact on the thermal surface distribution. Top: Temperature maps at the base of the envelope. Bottom: Magnetic field lines, with colors indicating field intensity (here not evolved). Simulations assumed a NS mass of $1.4M_{\odot}$ with the SLy4 EoS (see Fig. 1). Figure courtesy of Ascenzi et al (2024).
@@
Core-threading configurations are less well understood due to the complex physics within the inner core of NSs. Two main distinctions exist between crustal-confined and core-threading configurations: first, the field curvature of large-scale components differs by about an order of magnitude, corresponding to the star’s size versus the crust’s thickness; second, with the two regions having significantly different conductivities (see Fig. 8), the location of currents determines where Ohmic dissipation occurs and hence the timescale. As we have already observed in Fig. 21, the weaker Joule heating effect leads to core-threading models cooling much faster after the neutrino cooling era ($t\gtrsim 10^{4}$ yr). Thus, the observational appearance of a bright magnetar at late times hints for a consistent amount of electric currents in the crust. Note that the latter depends not only on how much the magnetic field penetrates in the core, but also on the complexity of the initial magnetic field: strong electrical currents can circulate for configurations like the ones discussed above, (e.g. Dehman et al 2023b), extended or not to the core.
@@
Conversely, for lower field strengths (bottom panels of Fig. 21), crustal-confined and core-threading models show similar behavior with minimal differences, due to the little relevance of magnetic effects (Ohmic heating and transport anisotropy). This has important observational implications (Marino et al, 2024) that we discuss in the next subsection, where we compare observational data to different models.
@@
Figure 25: Evolution of a mixed poloidal–toroidal core field under ambipolar diffusion. From left to right, columns show (a) Configuration of the magnetic field, where lines represent the poloidal magnetic field (labeled by the magnitude of $\alpha$) and colors the toroidal potential $\beta$; (b) and (c) density perturbations $\delta n_{n}$ and $\delta n_{c}$, respectively, both normalized to $n_{c0}$; (d) and (e) poloidal component of the neutron velocity, $\boldsymbol{v}n$, and ambipolar diffusion velocity, $\boldsymbol{v}{ad}$, where arrows represent the direction and colors the magnitude normalized to $R/t_{0}$. Rows correspond to different times: $t=0,\,t_{\zeta p},\,t_{\zeta g},\,t_{\zeta B}$, and $t_{ad}$. Figure courtesy of Castillo et al (2020).
@@
Prior to concluding this section, some more remarks on ambipolar diffusion in the core are in order. Despite some limitations, 2D (Castillo et al, 2017; Passamonti et al, 2017; Castillo et al, 2020; Viganò et al, 2021; Castillo et al, 2025; Moraga et al, 2025) and 3D (Igoshev and Hollerbach, 2023) studies demonstrate that, under some circumstances, ambipolar diffusion can drive the long-term reorganization and decay of magnetic fields in NS cores. As an illustrative example, Fig. 25 shows the results of a two-fluid simulation of an initial mixed large-scale-only poloidal–toroidal core field under ambipolar diffusion, assuming axial symmetry. The figure summarizes some of the expected key features, occurring at different characteristic timescales, from shorter to longer: the timescale for propagation of sound waves $t_{\zeta p}$ (characteristic of p-modes), the timescale associated with the Brunt-Väisälä frequency $t_{\zeta g}$ (characteristic of g-modes), the Alfvén crossing time, $t_{\zeta B}$ , and the ambipolar diffusion timescale $t_{ad}$, given by
@@
(see 25 for detailed definitions of the rest of characteristic timescales).
The simulation shows how the interplay of magnetic field and the two-fluid (neutral and charged fluids) dynamics drives the NS core through the following sequence of quasi-equilibria. Following a brief relaxation phase (where $t_{\zeta p}$, $t_{\zeta g}$, and $t_{\zeta B}$ represent rapid dynamical timescales on the order of fractions of a second), the system exhibits small density perturbations in its two components and establishes a nearly steady velocity field, approaching a twisted-torus equilibrium where magnetic, pressure, and buoyancy forces are almost balanced. This quasi-equilibrium is non-barotropic, as neutrons and charged particles contribute differently to the force balance (Castillo et al, 2020). By $t\approx t_{\zeta B}$, the toroidal magnetic field has been fully restructured, primarily persisting within closed poloidal loops. Density perturbations transition from being correlated with the magnetic field to uncorrelated, and the poloidal force imbalance diminishes, as evidenced by the reduced amplitude of velocities. From this point, ambipolar diffusion, that operates on a much longer timescale becomes the driver. From $t_{\zeta B}\to t_{ad}$, the magnetic force pushes charged particles relative to neutrons, transporting flux and reducing $|\delta n_{n}|$ while $|\delta n_{c}|$ grows until charged-particle gradients alone balance the field (note the scales at the basis of each panel). The last row of the figure shows both signatures: small $|\boldsymbol{v}_{ad}|$ compared with earlier times and suppressed neutron perturbations, consistent with a transition toward Grad-Shafranov like equilibria supported mainly by the charged fluid. Although this evolution is slow, it occurs more rapidly than in scenarios with stationary neutrons.
However, significant uncertainties remain. Realistic cores are expected to be superfluid and superconducting, which should at the very least alter the couplings between superfluid neutrons and superconducting protons (thus modifying the time-scales for the processes simulated in these studies) and even more importantly, modify the governing induction equation Glampedakis et al (2011); Graber et al (2015); Kantor and Gusakov (2018). We anticipate that this would be one very active line of research in the incoming years.
@@ 6.4 Comparison with observations
Magneto-thermal simulations of isolated NSs are particularly valuable as they enable direct comparison with observational data, notably quiescent thermal X-ray luminosities, as well as timing properties $P$ and $\dot{P}$. A key example is the work of Viganò et al (2013), who consistently re-analysed data from 40 isolated, thermally emitting NSs and showed that their phenomenological diversity can be explained by varying only the initial magnetic field, NS mass, and envelope composition. More recently, Potekhin et al (2020) conducted a complementary survey, presenting estimated ages, surface temperatures, and thermal luminosities of middle-aged NSs with relatively weak to moderately strong magnetic fields. Their comparison with theory demonstrated that the agreement between observational data and theoretical cooling curves improves significantly when models assume weak neutron superfluidity in the stellar core.
@@
In recent years, increasing attention has turned to the fast cooling scenario (Mendes et al, 2022; Marino et al, 2024). This interest was reinforced by the careful monitoring of three sources with well-determined ages that appear significantly colder by nearly an order of magnitude than other objects of comparable youth. These include two standard radio pulsars, PSR J0205++6449 ($P=70$ ms, $B_{p}=7\times 10^{12}$ G, age =841=841 yr; Kothes 2013) and PSR B2334++61 ($P=490$ ms, $B_{p}\sim 2\times 10^{13}$ G, age $\sim 7700$ yr; Yar-Uyaniker et al 2004), and one CCO CXOU J0852-4617 (age $\sim 2500$-5000 yr; Allen et al 2015).
@@
The secular cooling of NSs is influenced by the EoS, mass, magnetic field, and composition of the envelope, with the last three factors varying from star to star. By measuring the surface temperatures of numerous objects across a wide age range, NS cooling models (and consequently, the EoS) can be effectively constrained (Page et al, 2004). Cooling curves are generally categorized into standard (minimal) and enhanced regimes. The surface temperatures of observed neutron stars typically align with standard cooling models (Page et al, 2004; Potekhin et al, 2015b), although evidence of enhanced cooling has been noted for decades, with the Vela pulsar serving as a prime example. However, uncertainties in spectral energy distributions, precise ages, and accurate distances have hindered robust constraints on the equation of state (EoS) in this context.
@@
A detailed study of the three above-mentioned exceptionally cold, young, and nearby NSs was presented in Marino et al (2024). Reconciling theoretical models with these observations requires the inclusion of enhanced cooling processes, which in turn provides constraints on the NS EoS. A large set of simulations explored three representative EoSs spanning different cooling channels: SLy4 (Douchin and Haensel, 2001), which forbids enhanced cooling; BSK24 (Pearson et al, 2018); and GM1A (Gusakov et al, 2014), allowing for fast cooling. Simulations covered three masses (1.4, 1.6, and 1.8 M) with moderate magnetic fields ($\lesssim 7\times 10^{13}$ G) and an iron envelope, to prevent high luminosities that would be incompatible with these sources (see Fig. 21). The results, shown in Fig. 26, show that several scenarios fail to reproduce the faint thermal luminosities of the three cold sources. In particular, with the SLy4 EoS (orange curves), the sharp luminosity drop cannot be obtained for any mass or magnetic field configuration. By contrast, in the GM1A case with hyperons (blue/green curves), cooling can proceed rapidly enough to match the data. Similarly, for the BSK24 EoS, massive stars ($M\geq 1.6$ M) activate nucleon direct Urca, producing enhanced cooling tracks consistent with the observations. These results provide compelling evidence of enhanced cooling, showing that only EoSs (and compositions) allowing fast cooling within the first few thousand years can reproduce the observed thermal emission from this sample (Marino et al, 2024). It also reinforces the early suggestion by Aguilera et al (2008b) that DUrca cooling may be masked by strong magnetic fields in other NSs, potentially leading to misidentification. Importantly, the EoS should explain both exceptionally bright objects, such as magnetars, and extremely faint sources at young ages.
@@
Figure 26: Figure adapted from Marino et al (2024). Comparison between observational data and theoretical cooling curves. Standard rotation-powered pulsars are shown as squares and CCOs as circles. The 81 theoretical cooling curves used in our analysis are shown for three EoSs: SLy4 (orange), BSk24 (violet), and GM1A (blue). We consider three masses: $1.4\,M_{\odot}$ (dots), $1.6\,M_{\odot}$ (dashed), and $1.8\,M_{\odot}$ (solid). We explore nine initial polar surface dipolar fields from $B_{p}=1\times 10^{12}$ to $7\times 10^{13}$ G. For comparison only, we also plot three gray curves for stronger fields with an initial surface polar value of $10^{14}$, $3\times 10^{14}$, and $10^{15}$ G (BSk24, $1.6\,M_{\odot}$). In all cases, fields are crust-confined and initially purely large scale.
@@
Finally, we turn our attention to the rotational evolution of NSs. In Fig. 27 we present evolutionary tracks in the $P-\dot{P}$ diagram for a NS of 1.6 $M_{\odot}$ with varying initial magnetic field strengths, using a crustal-confined magnetic field configuration. Thin gray lines represent trajectories without field evolution, assuming a constant magnetic field, and appearing as straight lines in the diagram. In contrast, solid lines incorporating realistic field evolution deviate significantly from these models. Initially, the tracks coincide, as $B_{p}$ retains its initial value for early times ($t\lesssim 10^{3}-10^{5}$ yr). Over time, the field dissipates faster than the spin period evolves, causing the tracks to bend downward at nearly constant $P$. This behavior is suggested as the primary cause of the observed period clustering in isolated X-ray pulsars (Pons et al, 2013). The limiting period depends mainly on the initial magnetic field and crust-core interface resistivity. The significant differences between constant and realistic magnetic field models highlight the need to account for the interplay between temperature, magnetic, and rotational evolution.

### v2

@@ 6.2 Influence of boundary conditions on the long-term evolution
Thermal boundary conditions play a critical role in determining cooling timescales, as they govern energy losses through surface photon emission. Notably, significant differences arise when comparing non-magnetized envelope models (Gudmundsson et al., 1983; Potekhin et al., 1997) with magnetized ones (Potekhin et al., 2003, 2015b), for both light-element (hydrogen) and heavy-element (iron-like) compositions. Light-element envelopes, commonly used for accreting sources, produce luminosities up to an order of magnitude higher than heavy-element envelopes during the neutrino-cooling phase. Due to the efficient surface photon losses, light-element envelope models cool down faster once the photon-cooling era begins, rendering the objects hardly detectable in terms of thermal $X$-rays ($L_{X}\lesssim 10^{32}$ erg/s), much before than the heavy-element models. This effect is further amplified by strong magnetic fields, as first shown by Page and Sarmiento (1996) and more recently confirmed by Dehman et al. (2023a). They showed that the predicted thermal X-ray luminosity varies significantly based on multiple factors, particularly the envelope’s properties (see Sect. 2.2.5 for details). Their findings are summarized in Fig. 21, which compares outcomes under different assumptions about the magnetic field, composition (iron versus light elements like hydrogen, typical in accreted envelopes), or the internal current distribution (core-threading field versus crustal-confined field).
@@
Figure 21: Luminosity curves for four envelope models: non-magnetised heavy envelopes (solid, Gudmundsson et al. 1983), non-magnetised light envelopes (dots, Potekhin et al. 1997), magnetised light envelope (dashes, Potekhin et al. 2003), and magnetised heavy envelope (dot-dashes, Potekhin et al. 2015b). Left panels show models with crust-confined magnetic fields, while right panels show those with core-dominated magnetic fields, in all cases consisting of initial large-scale components only. Results are presented for two initial polar field strengths: $B=5\times 10^{14}$ G (top panels) and $B=10^{13}$ G (bottom panels). Image reproduced with permission from Dehman et al. (2023a), copyright by the authors.
@@
Beyond initial conditions, early-time ($t\lesssim 100$ yr) field reshaping and boundary-condition effects, the key question already mentioned in the previous section is arguably the location of the electric currents sustaining the magnetic field. Crustal-confined fields have been extensively studied and many works (Pons and Geppert, 2007; Viganò et al., 2012, 2013; Gourgouliatos et al., 2013; Gourgouliatos and Cumming, 2014b, a; Viganò et al., 2021; Gourgouliatos and Pons, 2019; De Grandis et al., 2020, 2022; Dehman et al., 2023c, b) generally agree on the overall picture of the Hall-driven dynamics in these configurations.
For typical field strengths of $10^{14}$ G, and starting from a predominantly poloidal dipolar field, we observe a stage dominated by the Hall drift (readjusting from initial conditions), which creates higher-order multipoles, followed by a quasi-stationary Ohmic stage. This structure, which has been called the Hall attractor (Gourgouliatos and Cumming, 2014a; Bransgrove et al., 2018), is characterized by a nearly constant angular velocity of the ”electron” fluid ($\Omega\approx j/en_{e}r$) along each poloidal field line, and proportional to the magnetic flux. It is worth noting that Hall drift can significantly accelerate magnetic field dissipation by steadily channeling magnetic energy to smaller scales, where Ohmic dissipation is more efficient.
@@
As an example, in Fig. 22 we show three snapshots of the evolution of a simple crustal-confined axisymmetric model, initially a $\ell=1$ poloidal field with $B_{p}=10^{14}$ G (labeled as model A14 in Viganò et al. 2013). Such very simple initial configuration allows one to analyse and capture several general basic features characterizing the Hall-dominated dynamics. Let us recap the most important facts:
@@
    At this stage, most of the current flows near the crust/core interface, where magnetic energy dissipation is governed by the resistivity of this region. In this particular model, a highly resistive layer in the nuclear pasta region causes rapid magnetic field decay, directly affecting the observable rotational properties of X-ray pulsars (Pons et al., 2013).
@@
Figure 23: Magnetic energy in the crust (normalized to the initial value) as a function of time, for the same model of Fig. 22. The solid lines correspond respectively to the total magnetic energy (black), the energy in the poloidal component (red), and the energy in the toroidal component (blue). The dashed line shows the evolution of the same model when the Hall term is deactivated (only Ohmic dissipation). Image reproduced with permission from Viganò et al. (2013), copyright by the authors.
@@
In 3D, details become even more relevant. Fig. 24 highlights the role of complex, multipolar magnetic structures close to the star surface in producing anisotropies in the temperature evolution. Such anisotropies have direct implications for the observable surface emission, potentially biasing cooling-age estimates and the interpretation of $X$-ray spectra, if oversimplified. This underscores the need for careful multi-dimensional treatments of the heat diffusion equation, rather than relying on one-dimensional cooling models (De Grandis et al., 2021; Igoshev et al., 2021, 2023; Dehman et al., 2023b; Ascenzi et al., 2024).
@@
Figure 24: 3D visualizations of different magnetic field configurations and their impact on the thermal surface distribution. Top: Temperature maps at the base of the envelope. Bottom: Magnetic field lines, with colors indicating field intensity (here not evolved). Simulations assumed a NS mass of $1.4M_{\odot}$ with the SLy4 EoS (see Fig. 1). Image reproduced with permission from Ascenzi et al. (2024), copyright by the author(s).
@@
Core-threading configurations are less well understood due to the complex physics within the inner core of NSs. Two main distinctions exist between crustal-confined and core-threading configurations: first, the field curvature of large-scale components differs by about an order of magnitude, corresponding to the star’s size versus the crust’s thickness; second, with the two regions having significantly different conductivities (see Fig. 8), the location of currents determines where Ohmic dissipation occurs and hence the timescale. As we have already observed in Fig. 21, the weaker Joule heating effect leads to core-threading models cooling much faster after the neutrino cooling era ($t\gtrsim 10^{4}$ yr). Thus, the observational appearance of a bright magnetar at late times hints for a consistent amount of electric currents in the crust. Note that the latter depends not only on how much the magnetic field penetrates in the core, but also on the complexity of the initial magnetic field: strong electrical currents can circulate for configurations like the ones discussed above, (e.g. Dehman et al. 2023b), extended or not to the core.
@@
Conversely, for lower field strengths (bottom panels of Fig. 21), crustal-confined and core-threading models show similar behavior with minimal differences, due to the little relevance of magnetic effects (Ohmic heating and transport anisotropy). This has important observational implications (Marino et al., 2024) that we discuss in the next subsection, where we compare observational data to different models.
@@
Figure 25: Evolution of a mixed poloidal–toroidal core field under ambipolar diffusion. From left to right, columns show (a) Configuration of the magnetic field, where lines represent the poloidal magnetic field and colors the toroidal potential; (b) and (c) density perturbations $\delta n_{n}$ and $\delta n_{c}$, respectively, both normalized to $n_{c0}$; (d) and (e) poloidal component of the neutron velocity, $\boldsymbol{v}_{n}$, and ambipolar diffusion velocity, $\boldsymbol{v}_{ad}$, where arrows represent the direction and colors the magnitude normalized to $R/t_{0}$. Rows correspond to different times: $t=0,\,t_{\zeta p},\,t_{\zeta g},\,t_{\zeta B}$, and $t_{ad}$. Image reproduced with permission from Castillo et al. (2020), copyright by the authors.
@@
Prior to concluding this section, some more remarks on ambipolar diffusion in the core are in order. Despite some limitations, 2D (Castillo et al., 2017; Passamonti et al., 2017; Castillo et al., 2020; Viganò et al., 2021; Castillo et al., 2025; Moraga et al., 2025) and 3D (Igoshev and Hollerbach, 2023) studies demonstrate that, under some circumstances, ambipolar diffusion can drive the long-term reorganization and decay of magnetic fields in NS cores. As an illustrative example, Fig. 25 shows the results of a two-fluid simulation of an initial mixed large-scale-only poloidal–toroidal core field under ambipolar diffusion, assuming axial symmetry. The figure summarizes some of the expected key features, occurring at different characteristic timescales, from shorter to longer: the timescale for propagation of sound waves $t_{\zeta p}$ (characteristic of p-modes), the timescale associated with the Brunt-Väisälä frequency $t_{\zeta g}$ (characteristic of g-modes), the Alfvén crossing time, $t_{\zeta B}$ , and the ambipolar diffusion timescale $t_{ad}$, given by
@@
(see Fig. 25 for detailed definitions of the rest of characteristic timescales).
The simulation shows how the interplay of magnetic field and the two-fluid (neutral and charged fluids) dynamics drives the NS core through the following sequence of quasi-equilibria. Following a brief relaxation phase (where $t_{\zeta p}$, $t_{\zeta g}$, and $t_{\zeta B}$ represent rapid dynamical timescales on the order of fractions of a second), the system exhibits small density perturbations in its two components and establishes a nearly steady velocity field, approaching a twisted-torus equilibrium where magnetic, pressure, and buoyancy forces are almost balanced. This quasi-equilibrium is non-barotropic, as neutrons and charged particles contribute differently to the force balance (Castillo et al., 2020). By $t\approx t_{\zeta B}$, the toroidal magnetic field has been fully restructured, primarily persisting within closed poloidal loops. Density perturbations transition from being correlated with the magnetic field to uncorrelated, and the poloidal force imbalance diminishes, as evidenced by the reduced amplitude of velocities. From this point, ambipolar diffusion, that operates on a much longer timescale becomes the driver. From $t_{\zeta B}\to t_{ad}$, the magnetic force pushes charged particles relative to neutrons, transporting flux and reducing $|\delta n_{n}|$ while $|\delta n_{c}|$ grows until charged-particle gradients alone balance the field (note the scales at the basis of each panel). The last row of the figure shows both signatures: small $|\boldsymbol{v}_{ad}|$ compared with earlier times and suppressed neutron perturbations, consistent with a transition toward Grad–Shafranov like equilibria supported mainly by the charged fluid. Although this evolution is slow, it occurs more rapidly than in scenarios with stationary neutrons.
However, significant uncertainties remain. Realistic cores are expected to be superfluid and superconducting, which should at the very least alter the couplings between superfluid neutrons and superconducting protons (thus modifying the time-scales for the processes simulated in these studies) and even more importantly, modify the governing induction equation (Glampedakis et al., 2011; Graber et al., 2015; Kantor and Gusakov, 2018). We anticipate that this would be one very active line of research in the incoming years.
@@ 6.4 Comparison with observations
Magneto-thermal simulations of isolated NSs are particularly valuable as they enable direct comparison with observational data, notably quiescent thermal X-ray luminosities, as well as timing properties $P$ and $\dot{P}$. A key example is the work of Viganò et al. (2013), who consistently re-analysed data from 40 isolated, thermally emitting NSs and showed that their phenomenological diversity can be explained by varying only the initial magnetic field, NS mass, and envelope composition. More recently, Potekhin et al. (2020) conducted a complementary survey, presenting estimated ages, surface temperatures, and thermal luminosities of middle-aged NSs with relatively weak to moderately strong magnetic fields. Their comparison with theory demonstrated that the agreement between observational data and theoretical cooling curves improves significantly when models assume weak neutron superfluidity in the stellar core.
@@
In recent years, increasing attention has turned to the fast cooling scenario (Mendes et al., 2022; Marino et al., 2024). This interest was reinforced by the careful monitoring of three sources with well-determined ages that appear significantly colder by nearly an order of magnitude than other objects of comparable youth. These include two standard radio pulsars, PSR J0205++6449 ($P=70$ ms, $B_{p}=7\times 10^{12}$ G, age =841=841 yr; Kothes 2013) and PSR B2334++61 ($P=490$ ms, $B_{p}\sim 2\times 10^{13}$ G, age $\sim 7700$ yr; Yar-Uyaniker et al. 2004), and one CCO CXOU J0852-4617 (age $\sim 2500$-5000 yr; Allen et al. 2015).
@@
The secular cooling of NSs is influenced by the EoS, mass, magnetic field, and composition of the envelope, with the last three factors varying from star to star. By measuring the surface temperatures of numerous objects across a wide age range, NS cooling models (and consequently, the EoS) can be effectively constrained (Page et al., 2004). Cooling curves are generally categorized into standard (minimal) and enhanced regimes. The surface temperatures of observed neutron stars typically align with standard cooling models (Page et al., 2004; Potekhin et al., 2015b), although evidence of enhanced cooling has been noted for decades, with the Vela pulsar serving as a prime example. However, uncertainties in spectral energy distributions, precise ages, and accurate distances have hindered robust constraints on the equation of state (EoS) in this context.
@@
A detailed study of the three above-mentioned exceptionally cold, young, and nearby NSs was presented in Marino et al. (2024). Reconciling theoretical models with these observations requires the inclusion of enhanced cooling processes, which in turn provides constraints on the NS EoS. A large set of simulations explored three representative EoSs spanning different cooling channels: SLy4 (Douchin and Haensel, 2001), which forbids enhanced cooling; BSK24 (Pearson et al., 2018); and GM1A (Gusakov et al., 2014), allowing for fast cooling. Simulations covered three masses (1.4, 1.6, and 1.8 M) with moderate magnetic fields ($\lesssim 7\times 10^{13}$ G) and an iron envelope, to prevent high luminosities that would be incompatible with these sources (see Fig. 21). The results, shown in Fig. 26, show that several scenarios fail to reproduce the faint thermal luminosities of the three cold sources. In particular, with the SLy4 EoS (orange curves), the sharp luminosity drop cannot be obtained for any mass or magnetic field configuration. By contrast, in the GM1A case with hyperons (blue/green curves), cooling can proceed rapidly enough to match the data. Similarly, for the BSK24 EoS, massive stars ($M\geq 1.6$ M) activate nucleon direct Urca, producing enhanced cooling tracks consistent with the observations. These results provide compelling evidence of enhanced cooling, showing that only EoSs (and compositions) allowing fast cooling within the first few thousand years can reproduce the observed thermal emission from this sample (Marino et al., 2024). It also reinforces the early suggestion by Aguilera et al. (2008b) that DUrca cooling may be masked by strong magnetic fields in other NSs, potentially leading to misidentification. Importantly, the EoS should explain both exceptionally bright objects, such as magnetars, and extremely faint sources at young ages.
@@
Figure 26: Comparison between observational data and theoretical cooling curves. Standard rotation-powered pulsars are shown as squares and CCOs as circles. The 81 theoretical cooling curves used in our analysis are shown for three EoSs: SLy4 (orange), BSk24 (violet), and GM1A (blue). We consider three masses: $1.4\,M_{\odot}$ (dots), $1.6\,M_{\odot}$ (dashed), and $1.8\,M_{\odot}$ (solid). We explore nine initial polar surface dipolar fields from $B_{p}=1\times 10^{12}$ to $7\times 10^{13}$ G. For comparison only, we also plot three gray curves for stronger fields with an initial surface polar value of $10^{14}$, $3\times 10^{14}$, and $10^{15}$ G (BSk24, $1.6\,M_{\odot}$). In all cases, fields are crust-confined and initially purely large scale. Image adapted from Marino et al. (2024).
@@
Finally, we turn our attention to the rotational evolution of NSs. In Fig. 27 we present evolutionary tracks in the $P-\dot{P}$ diagram for a NS of 1.6 $M_{\odot}$ with varying initial magnetic field strengths, using a crustal-confined magnetic field configuration. Thin gray lines represent trajectories without field evolution, assuming a constant magnetic field, and appearing as straight lines in the diagram. In contrast, solid lines incorporating realistic field evolution deviate significantly from these models. Initially, the tracks coincide, as $B_{p}$ retains its initial value for early times ($t\lesssim 10^{3}-10^{5}$ yr). Over time, the field dissipates faster than the spin period evolves, causing the tracks to bend downward at nearly constant $P$. This behavior is suggested as the primary cause of the observed period clustering in isolated X-ray pulsars (Pons et al., 2013). The limiting period depends mainly on the initial magnetic field and crust-core interface resistivity. The significant differences between constant and realistic magnetic field models highlight the need to account for the interplay between temperature, magnetic, and rotational evolution.